OpenStudy (anonymous):

1.Which equation represents the axis of symmetry of the function y = –2x² + 4x –6? y = 1 x = 1 x = 3 x = –3 2.What are the coordinates of the vertex of the graph of the function y = –x² + 6x –11? (3, –2) (3, 16) (–3, –29) (–3, –20) 3.What are the coordinates of the vertex of the graph of the function y = –3x² –12x + 3? (–2, 29) (2, –15) (2, –9) (–2, 15)

OpenStudy (zpupster):

x= -b/2a formula for axis of symmetry

OpenStudy (anonymous):

General form of equation: y=ax^2+bx+c To find axis of symmetry use discriminant x=(-b/2a) To find the coordinates of the vertex use the x from the discriminate and plug it into the function to find y

OpenStudy (anonymous):

thx guys but that really didnt help me any.. :/

OpenStudy (zpupster):

ax^2 + bx + c for 1) b=4, a= -2 axis of symmetry = -(4)/2(-2)

OpenStudy (anonymous):

1)Axis of symmetry: x=-(4)/2(2)=-4/4=-1; axis of symmetry is -1 2)Vertex: use the same equation to find x, x=-(6)/2(-1)=-6/-2=3 Then use that x in the original equation: y=-(3)^2+6(3)-11=-2 Vertex is at (3,-2)

OpenStudy (zpupster):

for 1) the a is -2 not 2, which changes the sign of your answer

OpenStudy (anonymous):

x=3?

OpenStudy (zpupster):

no 2 is ok no 3 is b

OpenStudy (zpupster):

you did it right differentiate set x to zero , get x and plug it back in to get y